OpenStudy (anonymous):

Does anyone know how to find the general solution of the following: (x)dy/dx-4y=exp(x)

sam (.sam.):

\[x\frac{dy}{dx}-4y=e^x\] Divide both sides by x \[\frac{dy}{dx}-\frac{4y}{x}=\frac{e^x}{x}\] We have integrating factor of -4/x \[\Large \mu(x)=e^{\int\limits -\frac{4}{x}}=e^{-4\ln(x)}=\frac{1}{x^4}\] Multiply all terms by 1/x^4 \[\frac{1}{x^4}\frac{dy}{dx}-\frac{4y}{x}\frac{1}{x^4}=\frac{e^x}{x^5}\] notice from LHS is a end-product rule, so \[\frac{d}{dx}(\frac{1}{x^4}y)=\frac{e^x}{x^5}\] Integrate both sides \[\frac{y}{x^4}=\int\limits \frac{e^x}{x^5} dx\] Solve

OpenStudy (anonymous):

Thank you so much :D

sam (.sam.):

you're welcome :)